# Space Group Equivariant Crystal Diffusion

### v1

@@ Abstract
Accelerating inverse design of crystalline materials with generative models has significant implications for a range of technologies. Unlike other atomic systems, 3D crystals are invariant to discrete groups of isometries called the space groups. Crucially, these space group symmetries are known to heavily influence materials properties. We propose SGEquiDiff, a crystal generative model which naturally handles space group constraints with space group invariant likelihoods. SGEquiDiff consists of an SE(3)-invariant, telescoping discrete sampler of crystal lattices; permutation-invariant, transformer-based autoregressive sampling of Wyckoff positions, elements, and numbers of symmetrically unique atoms; and space group equivariant diffusion of atomic coordinates. We show that space group equivariant vector fields automatically live in the tangent spaces of the Wyckoff positions. SGEquiDiff achieves state-of-the-art performance on standard benchmark datasets as assessed by quantitative proxy metrics and quantum mechanical calculations.
@@ 3 Preliminaries
Given a space group and a point $x\in\mathbb{R}^{3}$, the stabilizer group $G_{x}\equiv\{g|gx=x\}\subset G$ (also called the site symmetry group in materials science) is the finite subgroup of $G$ that leaves $x$ invariant. A Wyckoff position is then defined as the set of points with conjugate stabilizer groups, i.e., $\{x^{\prime}|\exists\ g\in G:G_{x^{\prime}}=gG_{x}g^{-1}\}$. Conceptually, if $g$ is a point group operation, this means that all points in a Wyckoff position are invariant to the same space group operations up to a change of basis. By convention, when $x$ is described with respect to the lattice basis $\{L_{1},L_{2},L_{3}\}$, the size of the orbit of $x$ in the unit cell, |$|\{gx|g\in G,gx\in[0,1)^{3}\}|$, is called the Wyckoff multiplicity. Wyckoff positions whose stabilizer groups are non-trivial, i.e., include more than the identity operation, are referred to as special Wyckoff positions as opposed to the general Wyckoff position defined by the identity stabilizer group. Wyckoff positions are labeled by multiplicity and Wyckoff letter, where the lexicographic ordering gives the Wyckoff positions in (partial) order of increasing multiplicity.
@@
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+g(t)d\mathbf{w},$ (1)
where $\mathbf{f}(\mathbf{x},t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $\mathbf{w}$ is the standard Wiener process, and $g(t)\in\mathbb{R}$ is the diffusion coefficient. Samples are generated by running the process backwards in time from $T$ to 0 via the reverse SDE,
$d\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})]dt+g(t)d\mathbf{\bar{w}},$ (2)
@@
See proof in Appendix A.1. The implication of this result is that a space group equivariant flow is automatically restricted to the manifold of its Wyckoff position, preventing the need for projection operations that lead to indeterminate probability distributions [39] or a discontinuous generation process [53, 76]. The result also implies that space group equivariant vector fields cannot transport atoms in higher-dimensional Wyckoff positions across lower-dimensional Wyckoff positions, motivating SDE-based sampling to overcome these barriers with Brownian motion. For example, in space group 192, we observed symmetrically unique regions of the 1D Wyckoff position 12k are split by the 0D Wyckoff position 4c (see Fig. 3).
@@ 4.1 Crystal representation
We represented $p(G)$ as a simple vector of logits and trained it with the cross-entropy loss. Now we describe our method for the remaining conditional distributions.
@@ 4.2 Telescoping discrete sampling of $L$
Besides physical constraints, lattice generation requires the flexibility to learn highly peaked distributions since small perturbations to a crystal lattice can significantly alter materials properties. In BaTiO3 for example, $0.03\mathrm{\AA}$ strain was found to increase the ferroelectric transition temperature by 500C and the remnant polarization by 250% [13]. In contrast, the range of conventional lattice lengths in the MP20 dataset is over 100Å. To address this challenge, we chose to discretize the lattice parameters to a resolution of 0.01Å. Naively, this resolution requires a softmax over $N_{l}=\mathcal{O}(10^{4})$ classes per lattice parameter to achieve a 100Å range. We overcame this poor scaling by telescoping the categorical distribution. See Figure 2 for a visual explanation. At a high level, the range of lattice parameters was first binned very coarsely, and a class $b_{1}$ was sampled from $p(b_{1}|l_{<i},G)$. Then, the selected class $b_{1}$ was further coarsely binned and one of these higher resolution bins was selected from $p(b_{2}|b_{1},l_{<i},G)$. This process was repeated $K$ times to achieve higher levels of resolution as
@@
where $l_{i}=b_{K}$ and $b_{0}=\emptyset$. Choosing $p(b_{k}|b_{k-1},l_{<i},G)$ to be a categorical distribution over a small number of classes $n_{l}\ll N_{l}$ achieves $(1/n_{l}^{K})=(1/N_{l})$ resolution with $\mathcal{O}(n_{l}K)\ll\mathcal{O}(N_{l})$ memory. We used $K=2$ and $n_{l}=100$ in our experiments. We represented conditioning on $l_{i}$ and $b_{k}$ by using min-max normalization and Gaussian random Fourier features [86]. Logits for $p(b_{k}|b_{k-1},l_{<i},G)$ were produced with an MLP.
To better align the distributions of partially complete lattices seen during training and inference, we employed noisy teacher forcing during training. Specifically, we minimized the negative log likelihood of ground truth lattice parameters conditioned on noisy lattice parameters. Lattice lengths and angles were augmented with uniform random noise with 0.3Å and 5 ranges, respectively.
4.3 Transformer-based sampling of $W$, $A$, and $n$
@@
We used attention masking to enforce unique token orderings per crystal. Specifically, atoms were ordered lexicographically first by Wyckoff letter and then by atomic number. The transformer input is invariant to permutations of any remaining ties. We additionally set logits of zero-dimensional Wyckoff positions which are already occupied to negative infinity to prevent sampling overlapping atoms at those locations. The transformer was fit with the cross-entropy loss. See section A.4 for details of our transformer architecture.
To aid generalization across the conditioning variables, we initialized embeddings from physical descriptors. For the 230 space groups, we formed initial one-hot features using the lattice centering type, crystal family, point group symbol, chirality, presence of inversion symmetry, and extra dimensions to differentiate collisions from screw and glide symmetries, yielding 62 features total. We built 231-dimensional Wyckoff position embeddings for the 1731 Wyckoff positions using one-hot features of Wyckoff multiplicities, Wyckoff dimensionalities, site symmetry symbols, space group features, and average Fourier features of the coordinates at vertices and centers of masses of the Wyckoff shapes. Element embeddings were initialized as those introduced by Xie and Grossman [96]. Similarly to the lattice sampler, min-max normalized lattice parameters were embedded with Gaussian random Fourier features. Learnable stop tokens were initialized from Gaussian noise.
@@ 4.4 Space group equivariant diffusion of $X$
Following Jiao et al. [38], we let the noise scale $\sigma_{t}$ follow the exponential scheduler: $\sigma_{0}=0$ and $\sigma_{t}=\sigma_{1}(\frac{\sigma_{T}}{\sigma_{1}})^{\frac{t-1}{T-1}}$ if $t>0$. As $\sigma_{t}$ gets large, the SGWN approaches the uniform distribution. Noisy samples $x_{t}$ were created by first sampling Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ and then reparameterizing as $x_{t}=x_{0}+\mathcal{P}_{x_{0}}(\sigma_{t}\epsilon)$ where $\mathcal{P}_{x_{0}}(\cdot)$ is an orthogonal projection to the tangent space of $x_{0}$’s Wyckoff shape.
For the backward process, we constructed the uniform prior $p(x_{T})$ on each Wyckoff position. To do so without redundantly assigning probability mass to symmetrically equivalent points, we placed uniform distributions on the relevant Wyckoff shapes. See Sec. A.3 for details. We used predictor-corrector sampling [82] to sample $x_{0}$.
@@
$L_{X}=\mathbb{E}_{x_{t}\sim p_{t}(x_{t}|x_{0})p(x_{0}),t\sim\mathcal{U}(0,T)}\big{[}||s_{\theta}(x_{t},t)-\lambda_{t}\nabla_{x_{t}}p(x_{t}|x_{0})||_{2}^{2}\big{]}$
where $\lambda_{t}=\mathbb{E}\big{[}||\nabla_{x_{t}}p(x_{t}|x_{0})||\big{]}^{-1}$ was employed for training stability. Approximate computation of $\lambda_{t}$ and $\nabla_{x_{t}}p(x_{t}|x_{0})$ is described in Sec. A.2.
@@ 5 Experiments
We evaluated SGEquiDiff on a suite of quantitative metrics and trained on two benchmark datasets: MP20 [97], which contains 45,231 experimentally known crystals with up to 20 atoms per unit cell, and the more challenging MPTS52 dataset [4], which contains 40,476 experimentally known crystals with up to 52 atoms per unit cell. Evaluations were conducted on a budget of 10,000 generated crystals. Following Xie et al. [97], we computed structural and compositional validity percentages using heuristics about interatomic distances and charge, respectively. We note that only $\sim$90of crystals in MP20 pass the compositional validity checker based on SMACT [16]. Of the 10,000 generated crystals, we randomly sampled 1,000 determined to be both structurally and compositionally valid. Of these 1,000 crystals, we determined how many were unique and novel (U.N.) with respect to the training data using pymatgen’s StructureMatcher [68] with stol=0.3, angle_tol=5, and ltol=0.2. U.N. crystals were used to compute (1) distribution distances between ground truth test and generated materials properties, including Wasserstein distances for atomic density $\rho$ and number of unique elements $N_{el}$ as well as Jensen-Shannon divergences (JSD) for space group $G$ and occupied Wyckoff dimensionalities $d_{\mathrm{Wyckoff}}$; (2) Central Moment Discrepancy (CMD) [103] up to 50 moments between ground truth test and generated crystal CrystalNN structural fingerprints [106]; and (3) structural and compositional diversity as measured by average pairwise $L_{2}$-distances between CrystalNN and Magpie [90] fingerprints, respectively. We also measured average sampling times per batch of 500 crystals on a single NVIDIA A40 GPU. We list hyperparameters and training times for SGEquiDiff in Sec. A.7.
Table 1: Results on the MP20 dataset. Best result is bolded, second best is underlined.
@@
Sampling time $\downarrow$ Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$ S.U.N. rate (%) $\uparrow$
(sec / batch) Structure Composition $\mathcal{D}^{\mathrm{MP20}}_{\mathrm{train}}$ $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{G}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition MP-2024
CDVAE [97] 906 99.99 85.66 98.2 0.6590 1.423 0.6957 0.4590 0.4821 0.6539 13.70 2.95
DiffCSP [38] 154 99.92 82.21 85.6 0.1454 0.4000 0.4638 0.2328 0.1766 0.9588 15.69 -
DiffCSP++ [39] 484 99.92 85.94 84.7 0.1658 0.5002 0.1608* 0.0449* 0.1079 0.9329 15.23 1.69
SymmCD [53] 139 88.24 86.76 87.7 0.1640 0.3213 0.1669* 0.0344 0.3233 0.9111 15.62 -
FlowMM (reported) [60] - 96.85 83.19 - - - - - - - - -
MatterGen [104] 3,689 100.0 82.6 86.3 0.2059 0.2416 0.4331 0.2129 0.1338 0.9889 15.67 7.77
SGEquiDiff 325 99.81 84.06 83.1 0.6247 0.1988 0.1769 0.0299 0.1918 0.9198 15.62 12.47
SGEquiDiff (T=1.1) 252 99.63 83.11 84.7 0.6603 0.2169 0.1748 0.0124 0.1965 0.8951 15.85 -
SGEquiDiff (T=1.5) 251 99.49 81.61 90.3 0.9310 0.3355 0.2055 0.0404 0.2830 0.8533 15.87 -
SGEquiDiff (T=2.0) 236 99.21 79.76 94.0 1.148 0.3887 0.2637 0.0620 0.3632 0.8304 16.26 -
SGEquiDiff (T=3.0) 262 98.95 77.63 96.8 2.167 0.5196 0.3298 0.0780 0.4538 0.7815 16.94 -
SGEquiDiff+LDiff 272 97.22 82.63 80.5 0.1794 0.1270 0.1617 0.0353 0.1686 0.9563 15.99 5.64
@@
* Uses fixed templates from the training data.
@@
Finally, for 100 random U.N. crystals, we conducted structure relaxations with expensive density functional theory (DFT) calculations and assessed whether the resulting crystals were thermodynamically stable. Aligned with analyses of DFT-calculated energies for experimentally observed crystals [85] and of DFT errors relative to experiment [84, 5], stability was defined as having a predicted energy above the hull less than 0.1 eV/atom with respect to crystals in the Materials Project v2024.11.14 [37]. Further details on the DFT calculations are in Sec. A.5. The fraction of stable materials was multiplied by the fraction of U.N. materials to calculate the stable, unique, and novel (S.U.N.) rate. We qualify that a low energy above the hull as predicted by DFT has been found to be a necessary but insufficient condition for synthesizability [85]; developing physical theories for predictive synthesis is an active area of research in the materials science community [58, 25, 41, 12, 66, 46]. Due to the computational expense of DFT calculations, we only computed S.U.N. rates on a subset of baseline models. While prior works pre-relax generated crystals with foundational machine learning interatomic potentials (MLIPs) before conducting DFT relaxations [60, 83, 104], we opted not to use MLIPs to avoid conflating their biases [17, 51] with our evaluations of the generative models.
@@
We compared our model to several prior methods. CDVAE [97] predicts lattices and numbers of atoms per unit cell with a VAE, subsequently sampling elements and coordinates with denoising diffusion. DiffCSP [38] jointly diffuses the lattice with the atoms and uses fractional coordinates in the atoms’ spatial representations. MatterGen [104] similarly diffuses the lattice, atom types, and atom positions, but leverages Cartesian atom coordinates. FlowMM [60] extends DiffCSP with the flow matching framework. DiffCSP++ [39] extends DiffCSP by enforcing space group constraints with projected diffusion of the lattice and atom coordinates; however, they do so without space group equivariant scores. SymmCD [53] enforces space group constraints with non-equivariant diffusion of atoms in the asymmetric unit followed by post-hoc projections to the Wyckoff positions.
As an ablation, we replaced our autoregressive lattice sampling with the space group constrained, DDPM-based [33] lattice sampling approach proposed in DiffCSP++ and used in SymmCD. We denote this variant of lattice sampling as SGEquiDiff+LDiff. We used the same cosine noise scheduler and hyperparameters as DiffCSP++, jointly diffusing the lattice and atom coordinates.

### v2

@@ Abstract
Accelerating inverse design of crystalline materials with generative models has significant implications for a range of technologies. Unlike other atomic systems, 3D crystals are invariant to discrete groups of isometries called the space groups. Crucially, these space group symmetries are known to heavily influence materials properties. We propose SGEquiDiff, a crystal generative model which naturally handles space group constraints with space group invariant likelihoods. SGEquiDiff consists of an SE(3)-invariant, telescoping discrete sampler of crystal lattices; permutation-invariant, transformer-based autoregressive sampling of Wyckoff positions, elements, and numbers of symmetrically unique atoms; and space group equivariant diffusion of atomic coordinates. We show that space group equivariant vector fields automatically live in the tangent spaces of the Wyckoff positions. SGEquiDiff achieves state-of-the-art performance on standard benchmark datasets as assessed by quantitative proxy metrics and quantum mechanical calculations. Our code is available at https://github.com/rees-c/sgequidiff.
@@ 3 Preliminaries
Given a space group and a point $x\in\mathbb{R}^{3}$, the stabilizer group $G_{x}\equiv\{g|gx=x\}\subset G$ (also called the site symmetry group in materials science) is the finite subgroup of $G$ that leaves $x$ invariant. A Wyckoff position is then defined as the set of points with conjugate stabilizer groups, i.e., $\{x^{\prime}|\exists\ g\in G:G_{x^{\prime}}=gG_{x}g^{-1}\}$. Conceptually, if $g$ is a point group operation, this means that all points in a Wyckoff position are invariant to the same space group operations up to a change of basis. By convention, when $x$ is described with respect to the lattice basis $\{L_{1},L_{2},L_{3}\}$, the size of the orbit of $x$ in the unit cell, $|\{gx|g\in G,gx\in[0,1)^{3}\}|$, is called the Wyckoff multiplicity. Wyckoff positions whose stabilizer groups are non-trivial, i.e., include more than the identity operation, are referred to as special Wyckoff positions as opposed to the general Wyckoff position defined by the identity stabilizer group. Wyckoff positions are labeled by multiplicity and Wyckoff letter, where the lexicographic ordering gives the Wyckoff positions in (partial) order of increasing multiplicity.
@@
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+\eta(t)d\mathbf{w},$ (1)
where $\mathbf{f}(\mathbf{x},t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $\mathbf{w}$ is the standard Wiener process, and $\eta(t)\in\mathbb{R}$ is the diffusion coefficient. Samples are generated by running the process backwards in time from $T$ to 0 via the reverse SDE,
$d\mathbf{x}=[\mathbf{f}(\mathbf{x},t)-\gamma(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})]dt+\eta(t)d\mathbf{\bar{w}},$ (2)
@@
See proof in Sec. A.1. This result implies that a space group equivariant flow is automatically restricted to the manifold of its Wyckoff position, preventing the need for projections that lead to indeterminate probability distributions [41] or a discontinuous generation process [55, 78]. The result also implies that if, for example, a 0D Wyckoff position (a point) lies in a 1D Wyckoff position (a line), then a space group equivariant vector field cannot move an atom along the line through the point. This is problematic because we observed examples where traversing between symmetrically unique regions of a 1D Wyckoff position requires moving through a 0D Wyckoff position (e.g., in space group 192, the 1D Wyckoff position 12k is split by the 0D Wyckoff position 4c. See Fig. 3). This issue motivates SDE-based sampling since Gaussian noise (projected onto the tangent space of the Wyckoff position) can stochastically perturb atoms through 0D Wyckoff positions while maintaining that the marginal distribution over atom coordinates is still space group invariant.
@@ 4.1 Crystal representation
Structuring the generation process in this way rather than jointly modeling all crystal components avoids tuning per-component loss weights [100], memory-intensive mask states [104], and post-hoc projections that are intractable to marginalize over [55]. We set $p(G)$ to the empirical training distribution. Now we describe our method for the remaining conditional distributions.
@@ 4.2 Telescoping discrete sampling of $L$
Besides physical constraints, lattice generation requires the flexibility to learn highly peaked distributions since small perturbations to a crystal lattice can significantly alter materials properties. In BaTiO3 for example, 0.03Å strain was found to increase the ferroelectric transition temperature by 500C and the remnant polarization by 250% [13]. In contrast, the range of conventional lattice lengths in the MP20 dataset is over 100Å. To address this challenge, we chose to discretize the lattice parameters to a resolution of 0.01Å. Naively, this resolution requires a softmax over $N_{l}=\mathcal{O}(10^{4})$ classes per lattice parameter to achieve a 100Å range. We overcame this poor scaling by telescoping the categorical distribution. See Figure 2 for a visual explanation. At a high level, the range of lattice parameters was first binned very coarsely, and a class $b_{1}$ was sampled from $p(b_{1}|l_{<i},G)$. Then, the selected class $b_{1}$ was further coarsely binned and one of these higher resolution bins was selected from $p(b_{2}|b_{1},l_{<i},G)$. This process was repeated $K$ times to achieve higher levels of resolution as
@@
where $l_{i}=b_{K}$ and $b_{0}=\emptyset$. Choosing $p(b_{k}|b_{k-1},l_{<i},G)$ to be a categorical distribution over a small number of classes $n_{l}\ll N_{l}$ achieves $(1/n_{l}^{K})=(1/N_{l})$ resolution with $\mathcal{O}(n_{l}K)\ll\mathcal{O}(N_{l})$ memory. We used $K=2$ and $n_{l}=100$ in our experiments. We represented conditioning on $l_{i}$ and $b_{k}$ by using min-max normalization and Gaussian random Fourier features [89]. Logits for $p(b_{k}|b_{k-1},l_{<i},G)$ were produced with an MLP. Space group constraints were enforced with hard-coded masking of lattice parameters and their logits.
To better align the distributions of partially complete lattices seen during training and inference, we employed noisy teacher forcing during training. Specifically, we minimized the negative log likelihood of the next ground truth lattice parameter conditioned on noisy versions of previous ground truth lattice parameters, where noise was masked to respect space group constraints. Lattice lengths and angles were augmented with uniform random noise with 0.3Å and 5 ranges, respectively. Noising was not applied at inference time.
4.3 Transformer-based sampling of $W$, AA, and $n$
@@
The number of atoms in the ASU $n$ is sampled implicitly via $p(\mathrm{stop}|w_{\leq i},a_{\leq i},L,G)$. We used attention masking to enforce unique token orderings per crystal during training and inference. Specifically, atoms were ordered lexicographically first by Wyckoff letter and then by atomic number. The transformer input is invariant to permutations of any remaining ties. We additionally set logits of zero-dimensional Wyckoff positions which are already occupied to negative infinity to prevent sampling overlapping atoms at those locations. The transformer was fit with the cross-entropy loss. See section A.4 for details of our transformer architecture. We did not use a positional encoding since the model does not need to distinguish between different orderings of the same atoms.
To aid generalization across the conditioning variables, we initialized embeddings from physical descriptors. Element embeddings were initialized as those introduced by Xie and Grossman [99]. Similarly to the lattice sampler, min-max normalized lattice parameters were embedded with Gaussian random Fourier features. Learnable stop tokens were initialized from Gaussian noise. We represented each of the 230 space groups with 62 initial features and each of the 1731 Wyckoff positions across space groups with 231 initial features. For details and an ablation against the symmetry features used in SymmCD [55], see Sec. A.5.1.
@@ 4.4 Space group equivariant diffusion of $X$
Following Jiao et al. [40], we set the noise scale $\sigma_{t}$ with the exponential scheduler: $\sigma_{0}=0$ and $\sigma_{t}=\sigma_{1}(\frac{\sigma_{T}}{\sigma_{1}})^{\frac{t-1}{T-1}}$ if $t>0$. As $\sigma_{t}$ gets large, the SGWN approaches the uniform distribution. Noisy samples $x_{t}$ were created by sampling Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ and reparameterizing as $x_{t}=x_{0}+\mathcal{P}_{x_{0}}(\sigma_{t}\epsilon)$ where $\mathcal{P}_{x_{0}}(\cdot)$ is an orthogonal projection to the tangent space of $x_{0}$’s Wyckoff shape. For the backward process, we constructed the uniform prior $p(x_{T})$ on each Wyckoff position. To do so without redundantly assigning probability mass to symmetrically equivalent points, we placed uniform distributions on the relevant Wyckoff shapes. See Sec. A.3 for details. We used predictor-corrector sampling [85] to sample $x_{0}$.
@@
$L_{X}=\mathbb{E}_{x_{t}\sim p_{t}(x_{t}|x_{0})p(x_{0}),t\sim\mathcal{U}(0,T)}\big[||s_{\theta}(x_{t},t)-\lambda_{t}\nabla_{x_{t}}\log p(x_{t}|x_{0})||_{2}^{2}\big]$
where $\lambda_{t}=\mathbb{E}\big[||\nabla_{x_{t}}\log p(x_{t}|x_{0})||\big]^{-1}$ was employed for training stability. Approximate computation of $\lambda_{t}$ and $\nabla_{x_{t}}\log p(x_{t}|x_{0})$ is described in Sec. A.2.
@@ 5 Experiments
We evaluated SGEquiDiff on two benchmark datasets: MP20 [100], containing 45,231 experimentally known crystals with up to 20 atoms per unit cell, and the more challenging MPTS52 dataset [4], containing 40,476 experimentally known crystals with up to 52 atoms per unit cell. Evaluations were conducted on 10,000 generated crystals. Following Xie et al. [100], we computed structural and compositional validity percentages using heuristics of interatomic distances and charge, respectively. We note that only $\sim$90% of crystals in MP20 pass the composition validity checker based on SMACT [16]. Of the 10,000 generated crystals, we randomly sampled 1,000 determined to be both structurally and compositionally valid. Of these 1,000 crystals, we determined how many were unique and novel with respect to the training data using pymatgen’s StructureMatcher [70] with stol=0.3, angle_tol=5, and ltol=0.2 (U.N. structures). We also followed SymmCD and calculated the fraction of 1,000 valid crystals with a unique and novel template (U.N. templates), where a template is defined as a space group and multiset of occupied Wyckoff positions. U.N. structures were used to compute (1) distribution distances between ground truth test and generated materials properties, including Wasserstein distances for atomic density $\rho$ and number of unique elements $N_{el}$ as well as Jensen-Shannon divergences (JSD) for space group $G$ and occupied Wyckoff dimensionalities $d_{\mathrm{Wyckoff}}$; (2) Central Moment Discrepancy (CMD) [106] up to 50 moments between ground truth test and generated crystal CrystalNN structural fingerprints [109]; and (3) structural and compositional diversity as measured by average pairwise $L_{2}$-distances between CrystalNN and Magpie [93] fingerprints, respectively. Average sampling times per batch of 500 crystals were measured on an NVIDIA A40 GPU. We list hyperparameters and training times for SGEquiDiff in Sec. A.8.
Table 1: MP-20 dataset. We report metrics of validity for 10,000 sampled crystals; uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for 100 valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. $T$ represents the inference-time temperature of discrete distributions in SGEquiDiff. Unless noted otherwise, TT=1.0. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
@@
Time $\downarrow$ Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$ S.U.N. (%) $\uparrow$
(s / batch) Structure Composition Template Structure $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{G}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition MP-2024
CDVAE [100] 906 99.99 85.66 15.5 98.2 0.6590 1.423 0.6957 0.4590 0.4821 0.6539 13.70 6.87
DiffCSP [40] 154 99.92 82.21 12.2 85.6 0.1454 0.4000 0.4638 0.2328 0.1766 0.9588 15.69 -
DiffCSP++ [41] 484 99.92 85.94 1.1 84.7 0.1658 0.5002 0.1608* 0.0449* 0.1079 0.9329 15.23 10.16
SymmCD [55] 139 88.24 86.76 9.6 87.7 0.1640 0.3213 0.1669 0.0344 0.3233 0.9111 15.62 -
[…]
MatterGen [107] 3,689 100.0 82.6 15.2 86.3 0.2059 0.2416 0.4331 0.2129 0.1338 0.9889 15.67 22.44
SGEquiDiff 226(12) 99.25(28) 86.16(32) 18.6(16) 80.0(16) 0.1938(1049) 0.2092(678) 0.1733(53) 0.0252(90) 0.1718(58) 0.9190(204) 15.64(7) 25.80
SGEquiDiff (TT=1.5) - 98.97(18) 83.36(52) 31.9(15) 90.1(14) 0.3810(1440) 0.3006(148) 0.2183(73) 0.0436(43) 0.2932(92) 0.8556(101) 15.82(18) -
SGEquiDiff (TT=2.0) - 98.40(37) 81.08(22) 42.8(31) 95.0(2) 0.6524(937) 0.3685(409) 0.2698(108) 0.0564(61) 0.3760(115) 0.8439(137) 16.28(17) -
SGEquiDiff (TT=3.0) - 97.84(53) 78.60(35) 57.1(32) 95.6(5) 1.473(267) 0.4962(177) 0.3367(204) 0.0751(41) 0.4656(132) 0.8177(101) 17.16(30) -
@@
* Uses fixed templates from the training data. $\dagger$ Values reported by Miller et al. [62].
@@
Finally, for 100 random U.N. structures, we conducted structure relaxations with expensive density functional theory (DFT) calculations and assessed whether the resulting crystals were thermodynamically stable. Aligned with analyses of DFT-calculated energies for experimentally observed crystals [88] and of DFT errors relative to experiment [87, 5], stability was defined as having a predicted energy above the hull less than 0.1 eV/atom with respect to crystals in the Materials Project v2022.8.23 [39]. Further details on the DFT calculations are in Sec. A.5. The fraction of stable materials was multiplied by the fraction of U.N. structures to calculate the stable, unique, and novel (S.U.N.) rate. We qualify that a low energy above the hull as predicted by DFT has been found to be a necessary but insufficient condition for synthesizability [88]; developing physical theories for predictive synthesis is an active area of research in the materials science community [60, 25, 43, 12, 68, 48]. Due to the computational expense of DFT calculations, we only computed S.U.N. rates on a subset of baseline models. While prior works pre-relax generated crystals with foundational machine learning interatomic potentials (MLIPs) before conducting DFT relaxations [62, 86, 107], we opted not to use MLIPs to avoid conflating their biases [17, 53] with our evaluations of the generative models.
@@
We compared our model to several prior methods. CDVAE (4.9M parameters) [100] predicts lattices and numbers of atoms per unit cell with a VAE and then samples elements and coordinates with denoising diffusion. DiffCSP (12.3M parameters) [40] jointly diffuses the lattice with the atoms using fractional atom coordinates. MatterGen (44.6M parameters) [107] similarly diffuses the lattice, atom types, and atom positions, but leverages Cartesian atom coordinates. FlowMM [62] extends DiffCSP with the flow matching framework. DiffCSP++ (12.3M parameters) [41] extends DiffCSP by enforcing space group constraints with projected diffusion of the lattice and atom coordinates; however, they do so without space group equivariant scores. SymmCD (60.4M parameters) [55] enforces space group constraints with non-equivariant diffusion of atoms in the asymmetric unit followed by post-hoc projections to the Wyckoff positions.
Table 2: MPTS-52 dataset. We report metrics of validity for 10,000 sampled crystals; uniqueness and novelty for 1,000 valid crystals; and diversity and distributional distances for valid, unique, and novel crystals. Best result is bolded, second best is underlined. For DiffCSP and DiffCSP++, we used the same diffusion corrector step sizes as for MP-20, and the DiffCSP++ batch size was decreased to 96 to avoid OOMs. CDVAE failed to train due to graph construction errors from isolated atoms. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
Time $\downarrow$ Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$
(s / batch) Structure Composition Template Structure $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{G}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition
DiffCSP [40] 467 67.47 55.8 19.8 79.8 1.189 0.5006 0.6900 0.0818 0.4836 0.8621 16.48
DiffCSP++ [41] 1230 99.87 77.52 1.0 86.9 0.8244 0.3692 0.2759 0.1154 0.3812 0.8457 15.69
SymmCD [55] 210 87.11 78.18 14.8 90.2 1.126 0.3506 0.2730 0.1146 0.4775 0.7843 15.36
SGEquiDiff 530(40) 97.79(37) 79.83(96) 38.7(25) 89.8(7) 0.6110(1592) 0.1736(428) 0.3104(62) 0.1300(142) 0.4186(28) 0.8613(292) 16.31(11)